OpenStudy (anonymous):

Solve for x: |2x + 12| = 18 x = 3, x = −3 x = 3, x = −15 x = 15, x = −15 x = −3, x = 15

OpenStudy (anonymous):

Hi :)

OpenStudy (anonymous):

you have two possible answers |2x + 12| = 18 so possibility 1 2x + 12 = 18 or 2x + 12 = -18 solve both equations for the two possibilities

OpenStudy (anonymous):